OpenStudy (anonymous):

hey how do i put (1+(1/x)/2-(3/x) in the form of a(x)/b(x)?

OpenStudy (anonymous):

\[\frac {1+\frac{1}{x}}{2-\frac{3}{x}}\]

OpenStudy (anonymous):

Is that what you mean?

OpenStudy (anonymous):

yes I'm still getting used to the typing on here

OpenStudy (anonymous):

Ok, It's already in the form a(x)/b(x)

OpenStudy (anonymous):

But you can multiply the numerator and denominator by x to get

OpenStudy (anonymous):

oh ok thanks man

OpenStudy (anonymous):

\[\frac{x+1}{2x-3}\]

OpenStudy (anonymous):

You're welcome.